OpenStudy (anonymous):

lim(h)/[(sqrt(2x+2h-5)) - (sqrt(2x-5))] h->0

OpenStudy (sleepyhead314):

Hello and Welcome to OpenStudy! :) do you mean something like: \[\lim_{h \rightarrow 0} (\sqrt{2x + 2h-5} - \sqrt{2x-5})\]

OpenStudy (anonymous):

yes but with h being a numerator

OpenStudy (aum):

\[ \lim_{h \rightarrow 0} \frac{h}{(\sqrt{2x + 2h-5} - \sqrt{2x-5})} \]

OpenStudy (sleepyhead314):

right, so first I put the denominator all under one squareroot so the denominator becomes sqrt(2x-2x-5+5+2h) which you can then add like terms and cancel out some stuff

OpenStudy (aum):

\[ \lim_{h \rightarrow 0} \frac{h}{(\sqrt{2x + 2h-5} - \sqrt{2x-5})} = \\ \lim_{h \rightarrow 0} \frac{h}{(\sqrt{2x + 2h-5} - \sqrt{2x-5})} * \frac{(\sqrt{2x + 2h-5} + \sqrt{2x-5})}{(\sqrt{2x + 2h-5} + \sqrt{2x-5})}= ?\\ \]

OpenStudy (sleepyhead314):

oh right, that's the way to do it ^_^"

OpenStudy (aum):

\[ \lim_{h \rightarrow 0} \frac{h}{(\sqrt{2x + 2h-5} - \sqrt{2x-5})} = \\ \lim_{h \rightarrow 0} \frac{h}{(\sqrt{2x + 2h-5} - \sqrt{2x-5})} * \frac{(\sqrt{2x + 2h-5} + \sqrt{2x-5})}{(\sqrt{2x + 2h-5} + \sqrt{2x-5})}= \\ \lim_{h \rightarrow 0} \frac{h(\sqrt{2x + 2h-5} + \sqrt{2x-5})}{2x+2h-5-2x+5} = \\ \lim_{h \rightarrow 0} \frac{h(\sqrt{2x + 2h-5} + \sqrt{2x-5})}{2h} = \\ \lim_{h\rightarrow 0} \frac{(\sqrt{2x + 2h-5} + \sqrt{2x-5})}{2} = \sqrt{2x-5}\\ \]

OpenStudy (anonymous):

Thanks, i didnt see the 2h at the bottom

OpenStudy (aum):

The h at the bottom cancels the h at the top left.